# Temperature Dependence of Heavy Quark Diffusion from (2+1)-flavor Lattice QCD

### v1

@@ 5.3.2 Quasi-particle model for charm and bottom degrees of freedom
As previously discussed, Langevin dynamics requires specifying the masses of the charm and bottom quarks. At high temperatures, these heavy quarks are expected to be the relevant carriers of heavy-flavor quantum numbers. This expectation has been confirmed in quenched QCD through lattice calculations of susceptibilities associated with vector and axial-vector currents, as well as the scalar density, in the deconfined phase [24]. These studies showed that the susceptibilities can be well described by a quasi-particle model with temperature-dependent quark masses. In full QCD with 2+1 flavors, lattice calculations of generalized charm susceptibilities further support this picture, indicating that for $T>195$ MeV, the dominant degrees of freedom are indeed charm quarks [72, 73, 74, 75]. Based on this, we determine the charm quark mass, the mean squared thermal velocity ${\langle\mathbf{v}^{2}\rangle}$, and the mean squared thermal momentum ${\langle\mathbf{p}^{2}\rangle}$ using a quasi-particle description. This methodology follows the previous HotQCD study in Ref. [21].
@@
$\chi_{c}(T)=\frac{6}{\pi}\int\frac{\mathrm{d}^{3}p}{(2\pi)^{3}}e^{-E_{p}/T},\ E_{p}=\sqrt{p^{2}+M_{c}^{2}(T)}\,.$ (5.8)
Similarly, for ${\langle\mathbf{v}^{2}\rangle}$ and ${\langle\mathbf{p}^{2}\rangle}$, we can write [24]:
${\langle\mathbf{v}^{2}\rangle}$ $=\frac{6}{\pi\chi_{c}(T)}\int\frac{\mathrm{d}^{3}p}{(2\pi)^{3}}\frac{p^{2}}{E_{p}^{2}}e^{-E_{p}/T}\,,$ (5.9)
${\langle\mathbf{p}^{2}\rangle}$ $=\frac{6}{\pi\chi_{c}(T)}\int\frac{\mathrm{d}^{3}p}{(2\pi)^{3}}p^{2}e^{-E_{p}/T}\,.$ (5.10)
In these equations, we have used Boltzmann distributions due to the large value of the charm-quark mass. To determine its temperature-dependent mass $M_{c}(T)$ for $T\geq 195$ MeV, we use the continuum-extrapolated lattice QCD results for the charm susceptibility $\chi_{c}(T)$ of the Wuppertal-Budapest Collaboration [76]. With these values of $M_{c}(T)$, we compute the mean squared thermal velocity ${\langle\mathbf{v}^{2}\rangle}$ and momentum ${\langle\mathbf{p}^{2}\rangle}$ using Eqs. (5.9) and (5.10), respectively.
A detailed analysis of generalized charm susceptibilities shows that for $T<195$ MeV, charm hadron-like excitations, such as charm mesons and baryons, become important [73, 74, 75]. In this regime, the total charm pressure can be approximated as the sum of partial pressures from charm mesons, charm baryons, and charm quarks [73, 74, 75]. Recent continuum-extrapolated results for these partial pressures have been estimated in Ref. [75]. Therefore, in the temperature range below $195$ MeV, it is more appropriate to estimate ${\langle\mathbf{v}^{2}\rangle}$ and ${\langle\mathbf{p}^{2}\rangle}$ using an extended quasi-particle model:
${\langle\mathbf{v}^{2}\rangle}=\frac{1}{\pi\chi_{c}(T)}\sum_{i}g_{i}\int\frac{\mathrm{d}^{3}p}{(2\pi)^{3}}\frac{p^{2}}{(E^{i}_{p})^{2}}e^{-E_{p}^{i}/T}\,,$ (5.11)
${\langle\mathbf{p}^{2}\rangle}=\frac{1}{\pi\chi_{c}(T)}\sum_{i}g_{i}\int\frac{\mathrm{d}^{3}p}{(2\pi)^{3}}p^{2}e^{-E_{p}^{i}/T}\,,$ (5.12)
@@
Unfortunately, not much is known about the open bottom degrees of freedom in (2+1)-flavor QCD from lattice QCD calculations. Based on the heavy-quark symmetry arguments, we assume that bottom degrees of freedom behave similar to the charm degrees of freedom. Namely, we assume that for $T\geq 195$ MeV, bottom quarks are the dominant degrees of freedom, while for lower temperature we also use the extended quasi-particle model that includes bottom quarks and open beauty mesons and baryons. Due to absence of lattice QCD data, we rather use a hadron resonance gas (HRG) model for the bottom susceptibility to determine the temperature dependent quasi-particle bottom mass $M_{b}(T)$ along with the values of ${\langle\mathbf{v}^{2}\rangle}$ and ${\langle\mathbf{p}^{2}\rangle}$. Again, as discussed in the Appendix C, the values of ${\langle\mathbf{v}^{2}\rangle}$ and ${\langle\mathbf{p}^{2}\rangle}$ for bottom degrees of freedom are not very sensitive to our assumptions, cf. Tab. 4.
@@ 5.3.3 The temperature dependence $\kappa$ and $D_{s}$ for charm and bottom quarks
Having determined the mean squared thermal velocity and momentum of charm and bottom quarks above the crossover temperature, we can now compute the total heavy-quark momentum and spatial diffusion coefficient. Where relevant, we will also compare these results to the infinite-mass limit. In Fig. 18, we present our final results for the total momentum diffusion coefficient $\kappa$, obtained using the thermal averages listed in Tabs. 3 and 4.
The uncertainties in $\kappa$ arising from variations in the thermal averages (due to different quasi-particle models) are much smaller than the dominant uncertainties associated with the determination of $\kappa_{E}$ and $\kappa_{B}$. Moreover, the dependence of $\kappa$ on the heavy-quark mass is quite small up to the highest temperatures for which the inverse mass expansion remains valid.
@@
Figure 19: Final results for the spatial diffusion coefficients $2\pi TD_{s}$, for the charm, bottom and infinitely-heavy quarks, compared with several other estimations and determinations. We include the QPMp estimates [77], the bayesian analysis of [78], the data-driven result of the ALICE collaboration and the result extracted from T-matrix formalism [79]. The dotted line represents the strongly coupled estimated based on AdS/CFT [28]. In addition, we also include the NLO perturbative QCD estimate [13], where the band corresponds to variations of the renormalization scale from $\mu=\pi T$ (upper boundary) to $\mu=4\pi T$ (lower boundary). The central value of the perturbative band is obtained using a scale $\mu=2\pi T$. We have split the figure into two parts: (right) one for the low temperature behavior and (left) a second one with the high temperature behavior, where we have only included the bottom and infinitely heavy quark for $T=1$ GeV.
@@
For $195$ MeV $\leq T\leq$ 293 MeV, our results agree with the previous HotQCD results [20, 21]. Additionally, our results show a significant increase in $2\pi TD_{s}$ with increasing temperature, and are in agreement with the NLO perturbative predictions for a wide range of temperatures. At low temperatures, say $T<300$ MeV this agreement is rather accidental, as the perturbative result cannot be trusted at such low temperature.
The heavy-quark diffusion coefficient has also been studied in the context of strongly coupled $N=4$ supersymmetric gauge theory using the AdS/CFT correspondence [28, 80]. These calculations give $D_{s}=2/(\pi T\sqrt{\lambda})$, where $\lambda=g^{2}N_{c}$ is the ’t Hooft coupling. By setting $N_{c}=3$, and identifying $g^{2}$ with the QCD coupling $g^{2}_{N_{f}=3}(\mu=2\pi T)$, we obtain a strong-coupling estimate for $D_{s}$ based on AdS/CFT, shown as the black dashed line in Fig. 19. Remarkably, our lattice results for $D_{s}$ at the four lowest temperatures are close to this prediction. In addition, we note that another strong-coupling approach in Ref. [81] also yields $2\pi TD_{s}\sim 1$ for $T<250$ MeV.
In Fig. 19, we compare our result on $D_{s}$ with other phenomenological determinations of the heavy-quark diffusion coefficient. Our result is systematically lower than those derived from Bayesian analysis of HIC data [78], and is also lower than the data-driven result of the ALICE collaboration at $T\sim 150$ MeV [82]. However, our results are in agreement with those obtained using the T-matrix formalism [79, 83]. In Fig. 19, we also show the latest results from the Catania group [77] obtained by using an refined version of their original quasi-particle description, known as QPMp. Their corresponding results are in agreement with our determination for bottom quark at low temperature, but remain much larger for charm quark and over-predict $D_{s}$ for both charm and bottom quarks for temperatures $T>250$ MeV.
@@ 6 Conclusions
In this paper, we present the first complete (2+1)-flavor lattice QCD calculation of the heavy-quark diffusion coefficient for a wide range of temperatures. For low temperatures, we extracted $\kappa$ at a almost physical point, using a physical strange-quark mass $m_{s}$ and near-physical light-quark mass $m_{l}=m_{s}/20$, corresponding to a pion mass of $m_{\pi}\simeq 160$ MeV. On the other hand, for high temperatures we use a physical strange-quark mass $m_{s}$ and $m_{l}=m_{s}/5$, corresponding to a pion mass of $m_{\pi}\simeq 320$ MeV because light quark mass effects are not important in this region. This study extends previous calculations of the HotQCD collaboration [20, 21] to the physical point for low temperatures, extending the temperature range to 163 MeV $\leq T\leq$ 10000 MeV.
Our analysis of the chromo-electric correlation function employs the correct high-energy behavior of the corresponding spectral function, computed in Refs. [54, 58]. Although this correction does not significantly change the extracted value of the heavy-quark diffusion coefficient within errors, it resolves the puzzle of the large rescaling factor $K_{E}$ observed in previous works [19, 20]. As discussed in Section 4.2, this improvement shows, for the first time, a direct agreement between lattice data and the perturbative NLO prediction at the level of the correlation function. This, in turn, reinforces the validity of our spectral function reconstruction strategy.
Our results show that close the crossover temperature, the spatial heavy-quark diffusion coefficient approaches the strongly coupled AdS/CFT estimate [28, 80], supporting the picture of the quark-gluon plasma as a strongly-coupled system and an almost perfect fluid, in which heavy quarks rapidly thermalize [4, 5, 6]. At high temperatures, our results are consistent within errors with the NLO perturbative QCD prediction [13]. Across the entire temperature range, the lattice QCD results remain systematically lower than the estimates from Bayesian reconstructions [78] and data-driven extractions [10], also showing a weaker dependence on the heavy-quark mass compared to certain phenomenological models.
@@
Acknowledgments
This material is based on work supported by the U.S. Department of Energy, Office of Science, Office of Nuclear Physics under Contract No. DE-SC0012704 and by the Scientific Discovery through Advanced Computing (SciDAC) award "Fundamental Nuclear Physics at the Exascale and Beyond" and the Topical Collaboration in Nuclear Theory "Heavy-Flavor Theory (HEFTY) for QCD Matter".
Also, L.A., O.K., S.S., and J.H.W. acknowledge support from the Deutsche Forschungsgemeinschaft (DFG) through the CRC-TR 211 "Strong-interaction matter under extreme conditions" (Project No. 315477589 - TRR 211). J.H.W.’s research is also funded by the DFG under project number 417533893/GRK2575 "Rethinking Quantum Field Theory". J.H.W. acknowledges the support by the State of Hesse within the Research Cluster ELEMENTS (Project ID 500/10.006).
This research used computing time provided by the ALCC and INCITE programs at the Oak Ridge Leadership Computing Facility, a DOE Office of Science User Facility supported under contract no. DE-AC05-00OR22725; the National Energy Research Scientific Computing Center (NERSC), a DOE Office of Science User Facility at Lawrence Berkeley National Laboratory under Contract No. DE-AC02-05CH11231; PRACE awards on JUWELS at GCS@FZJ, Germany; Marconi100 at CINECA, Italy; and the LUMI-G supercomputer at the Finnish IT Center for Science (CSC). Computations were also performed on facilities of the USQCD Collaboration funded by the DOE Office of Science, in particular on the 21-g cluster at the Thomas Jefferson National Accelerator Facility (JLAB). In addition, parts of this work were performed on the GPU cluster of the University of Bielefeld, supported by HPC.NRW.
@@ Appendix B Fits of the chromo-electric and chromo-magnetic correlators
Appendix C Quasi-particle model for charm and bottom agrees of freedom
@@
Figure 24: Left: Lattice results together with the interpolation of the charm susceptibility $\chi_{2}^{c}$ from the BW collaboration [76] with the recent $\chi_{4}^{c}$ results from the HotQCD collaboration [75]. For comparison, we also show the prediction from a full hadron resonance gas model QM-HRG. Right: interpolation of the charm-quark mass extracted from the susceptibility data of the BW and HotQCD collaborations. The orange point indicates the interpolated value at $T=188$ MeV.
@@
Figure 25: Charm and bottom partial pressures. Left: charm quark partial pressure obtained by HotQCD determination [75] compared to the charm meson, charm baryon and total charm hadron pressure from a 1S1P HRG model. Right: bottom partial pressures, showing results from the 1S1P HRG (dotted lines) and full HRG (dashed lines). These estimates are subsequently used to extract the bottom-quark mass as a function of temperature across the different models discussed in the text.
@@
$T$ [MeV] $M_{c_{1}}$[GeV] ${\langle\mathbf{v}^{2}\rangle}_{c_{1}}$ $\langle\mathbf{\hat{p}}^{2}\rangle_{c_{1}}$ $M_{c_{2}}$[GeV] ${\langle\mathbf{v}^{2}\rangle}_{c_{2}}$ $\langle\mathbf{\hat{p}}^{2}\rangle_{c_{2}}$
@@
Table 3: Mean squared thermal velocity ${\langle\mathbf{v}^{2}\rangle}$ and scaled mean squared thermal momentum $\langle\mathbf{\hat{p}}^{2}\rangle\equiv{\langle\mathbf{p}^{2}\rangle}/(3M_{c}T)$ of charm degrees of freedom extracted from the quasi-particle models discussed in the text. Here, $M_{c}$ is the charm-quark mass in the corresponding quasi-particle model.
$T$ [MeV] $M_{b_{1}}$[GeV] ${\langle\mathbf{v}^{2}\rangle}_{b_{1}}$ $\langle\mathbf{\hat{p}}_{b_{1}}^{2}\rangle$ $M_{b_{2}}$[GeV] ${\langle\mathbf{v}^{2}\rangle}_{b_{2}}$ $\langle\mathbf{\hat{p}}_{b_{2}}^{2}\rangle$ $M_{b_{3}}$[GeV] ${\langle\mathbf{v}^{2}\rangle}_{b_{3}}$ $\langle\mathbf{\hat{p}}_{b_{3}}^{2}\rangle$
163 5.1357 0.0852 1.1227 5.5290 0.0824 1.0752 —— —— ——
174 5.1040 0.0905 1.1364 5.4816 0.0883 1.0811 —— —— ——
188 5.0591 0.0972 1.1547 5.4198 0.0958 1.0889 —— —— ——
195 5.0591 0.1041 1.1141 —— —— —— 4.5000 0.1174 1.1117
220 5.0591 0.1168 1.1210 —— —— —— 4.5000 0.1308 1.1265
251 5.0591 0.1325 1.1285 —— —— —— 4.5000 0.1470 1.1450
293 5.0591 0.1519 1.1507 —— —— —— 4.5000 0.1682 1.1702
352 5.0591 0.1781 1.1824 —— —— —— 4.5000 0.1966 1.2062
400 5.0591 0.1984 1.2085 —— —— —— 4.5000 0.2186 1.2358
444 5.0591 0.2163 1.2326 —— —— —— 4.5000 0.2380 1.2632
500 5.0591 0.2383 1.2637 —— —— —— 4.5000 0.2615 1.2985
1000 5.0591 0.3982 1.5547 —— —— —— 4.5000 0.4299 1.6304
Table 4: Mean squared thermal velocity ${\langle\mathbf{v}^{2}\rangle}$ and scaled mean squared thermal momentum $\langle\mathbf{\hat{p}}^{2}\rangle\equiv{\langle\mathbf{p}^{2}\rangle}/(3M_{b}T)$ of bottom degrees of freedom extracted from the quasi-particle models discussed in the text. Here, $M_{b}$ is the bottom-quark mass in the corresponding quasi-particle model.
@@
As discussed in the main text, a quasi-particle description is necessary to estimate the heavy-quark diffusion coefficient beyond LO in the inverse heavy-quark mass. In this section of the appendix, we detail the quasi-particle model used to compute ${\langle\mathbf{v}^{2}\rangle}$ and ${\langle\mathbf{p}^{2}\rangle}$, based on available lattice QCD results. A model of this kind was first introduced in Ref. [24] to describe the temperature dependence of vector and axial-vector current susceptibilities for heavy (charm and bottom) quarks in quenched QCD. In Ref. [24] it was shown that ${\langle\mathbf{v}^{2}\rangle}$ obtained from the ratio of susceptibility of the vector current and the quark number susceptibility calculated on the lattice agrees with quasi-particle model.
Generalized charm susceptibilities have been computed throughout the years in 2+1-flavor lattice QCD [72, 74, 75, 76]. Below the crossover temperature, their behavior can be understood within the framework of a hadron resonance gas (HRG) model. In this approach, the particle content includes not only experimentally observed charm hadrons listed by the Particle Data Group, but also additional states predicted by quark models [72, 74, 75]. This extended version is known as the QM-HRG model.
Above the chiral crossover temperature, charm susceptibilities can be described using a model in which the total charm pressure is given by the sum of partial pressures from charm mesons, charm baryons, and charm quarks [73, 74, 75]:

### v2

@@ 5.3.2 Quasi-particle model for charm and bottom degrees of freedom
As previously discussed, Langevin dynamics requires specifying the masses of the charm and bottom quarks. At high temperatures, these heavy quarks are expected to be the relevant carriers of heavy-flavor quantum numbers. This expectation has been confirmed in quenched QCD through lattice calculations of susceptibilities associated with vector and axial-vector currents, as well as the scalar density, in the deconfined phase Petreczky:2008px . This study showed that the susceptibilities can be well described by a quasi-particle model with temperature-dependent quark masses Petreczky:2008px . In full QCD with 2+1 flavors, lattice calculations of generalized charm susceptibilities further support this picture, indicating that for $T>195$ MeV, the dominant degrees of freedom are indeed charm quarks Bazavov:2014yba ; Mukherjee:2015mxc ; Bazavov:2023xzm ; Kaczmarek:2025dqt . Based on this, we determine the charm quark mass, the mean squared thermal velocity $\langle\mathbf{v}^{2}\rangle$}, and the mean squared thermal momentum $\langle\mathbf{p}^{2}\rangle$} using a quasi-particle description. This methodology follows the previous HotQCD study in Ref. Altenkort:2023eav .
@@
$\chi_{c}(T)=\frac{6}{\pi T}\int\frac{\mathrm{d}^{3}p}{(2\pi)^{3}}e^{-E_{p}/T},\ E_{p}=\sqrt{p^{2}+M_{c}^{2}(T)}\,$, (5.9)
where $M_{c}(T)$ is the temperature dependent quasi-particle mass. Similarly, for $\langle\mathbf{v}^{2}\rangle$} and $\langle\mathbf{p}^{2}\rangle$}, we can write Petreczky:2008px :
${\langle\mathbf{v}^{2}\rangle}$ $=\frac{6}{\pi T\chi_{c}(T)}\int\frac{\mathrm{d}^{3}p}{(2\pi)^{3}}\frac{p^{2}}{E_{p}^{2}}e^{-E_{p}/T}\,$, (5.10)
${\langle\mathbf{p}^{2}\rangle}$ $=\frac{6}{\pi T\chi_{c}(T)}\int\frac{\mathrm{d}^{3}p}{(2\pi)^{3}}p^{2}e^{-E_{p}/T}\,.$ (5.11)
In these equations, we have used Boltzmann distributions due to the large value of the charm-quark mass. To determine the temperature-dependent mass $M_{c}(T)$ for $T\geq 195$ MeV, we use the continuum-extrapolated lattice QCD results for the charm susceptibility $\chi_{c}(T)$ of the Wuppertal-Budapest Collaboration Bellwied:2015lba . With these values of $M_{c}(T)$, we compute the mean squared thermal velocity $\langle\mathbf{v}^{2}\rangle$} and momentum $\langle\mathbf{p}^{2}\rangle$} using Eqs. (5.10) and (5.11), respectively.
A detailed analysis of generalized charm susceptibilities shows that for $T<195$ MeV, charm hadron-like excitations, such as charm mesons and baryons, become important Mukherjee:2015mxc ; Bazavov:2023xzm ; Kaczmarek:2025dqt . In this regime, the total charm pressure can be approximated as the sum of the partial pressures from charm mesons, charm baryons, and charm quarks Mukherjee:2015mxc ; Bazavov:2023xzm ; Kaczmarek:2025dqt . Recent continuum-extrapolated results for these partial pressures have been estimated in Ref. Kaczmarek:2025dqt . Therefore, in the temperature range below $195$ MeV, it is more appropriate to estimate $\langle\mathbf{v}^{2}\rangle$} and $\langle\mathbf{p}^{2}\rangle$} using an extended quasi-particle model:
${\langle\mathbf{v}^{2}\rangle}=\frac{1}{\pi T\chi_{c}(T)}\sum_{i}g_{i}\int\frac{\mathrm{d}^{3}p}{(2\pi)^{3}}\frac{p^{2}}{(E^{i}_{p})^{2}}e^{-E_{p}^{i}/T}\,$, (5.12)
${\langle\mathbf{p}^{2}\rangle}=\frac{1}{\pi T\chi_{c}(T)}\sum_{i}g_{i}\int\frac{\mathrm{d}^{3}p}{(2\pi)^{3}}p^{2}e^{-E_{p}^{i}/T}\,$, (5.13)
@@
Unfortunately, not much is known about the open bottom degrees of freedom in (2+1)-flavor QCD from lattice QCD calculations. Based on the heavy-quark symmetry arguments, we assume that bottom degrees of freedom behave similarly to the charm degrees of freedom. Namely, we assume that for $T\geq 195$ MeV, bottom quarks are the dominant degrees of freedom, while for lower temperature we also use the extended quasi-particle model that includes bottom quarks and open beauty mesons and baryons. Due to absence of lattice QCD data, we rather use a hadron resonance gas (HRG) model for the bottom susceptibility to determine the temperature dependent quasi-particle bottom mass $M_{b}(T)$ along with the values of $\langle\mathbf{v}^{2}\rangle$} and $\langle\mathbf{p}^{2}\rangle$}. Again, as discussed in the Appendix C, the values of $\langle\mathbf{v}^{2}\rangle$} and $\langle\mathbf{p}^{2}\rangle$} for bottom degrees of freedom are not very sensitive to our assumptions, cf. Tab. 4.
@@ 5.3.3 The temperature dependence $\kappa$ and $D_{s}$ for charm and bottom quarks
Having determined the mean squared thermal velocity and momentum of charm and bottom quarks above the crossover temperature, we can now compute the total heavy-quark momentum and spatial diffusion coefficient. Where relevant, we will also compare these results to the infinite-mass limit. In Fig. 18, we present our final results for the total momentum diffusion coefficient $\kappa$, obtained using the thermal averages listed in Tabs. 3 and 4. The uncertainties in $\kappa$ arising from variations in the thermal averages (due to different quasi-particle models) are much smaller than the dominant uncertainties associated with the determination of $\kappa_{E}$ and $\kappa_{B}$. Moreover, the dependence of $\kappa$ on the heavy-quark mass is quite small up to the highest temperatures for which the inverse mass expansion remains valid.
@@
Figure 19: Final results for the spatial diffusion coefficients $2\pi TD_{s}$, for the charm, bottom and infinitely-heavy quarks, compared with several other estimations and determinations. We include the QPMp estimates Sambataro:2024mkr , the bayesian analysis of Xu:2017obm , the data-driven result of the ALICE collaboration and the result extracted from T-matrix formalism Tang:2023tkm . The dotted line represents the strongly coupled estimated based on AdS/CFT Casalderrey-Solana:2006fio . In addition, we also include the NLO perturbative QCD estimate Caron-Huot:2007rwy , where the band corresponds to variations of the renormalization scale from $\mu=\pi T$ (upper boundary) to $\mu=4\pi T$ (lower boundary). The central value of the perturbative band is obtained using a scale $\mu=2\pi T$. We have split the figure into two parts: (right) one for the low temperature behavior and (left) a second one with the high temperature behavior, where we have only included the bottom and infinitely heavy quark for $T=1$ GeV.
@@
For $195$ MeV $\leq T\leq$ 293 MeV, our results agree with the previous HotQCD results Altenkort:2023oms ; Altenkort:2023eav . Additionally, our results show a significant increase in $2\pi TD_{s}$ with increasing temperature, and are in agreement with the NLO perturbative predictions for a wide range of temperatures. At low temperatures, say $T<300$ MeV this agreement is rather accidental, as the perturbative result cannot be trusted at such low temperature.
The heavy-quark diffusion coefficient has also been studied in the context of strongly coupled $N=4$ supersymmetric gauge theory using the AdS/CFT correspondence Casalderrey-Solana:2006fio ; Herzog:2006gh . These calculations give $D_{s}=2/(\pi T\sqrt{\lambda})$, where $\lambda=g^{2}N_{c}$ is the ’t Hooft coupling. By setting $N_{c}=3$, and identifying $g^{2}$ with the QCD coupling $g^{2}_{N_{f}=3}(\mu=2\pi T)$, we obtain a strong-coupling estimate for $D_{s}$ based on AdS/CFT, shown as the black dashed line in Fig. 19. Remarkably, our lattice results for $D_{s}$ at the four lowest temperatures are close to this prediction. In addition, we note that another strong-coupling approach in Ref. Andreev:2017bvr also yields $2\pi TD_{s}\sim 1$ for $T<250$ MeV.
In Fig. 19, we compare our result on $D_{s}$ with other phenomenological determinations of the heavy-quark diffusion coefficient. Our result is systematically lower than those derived from Bayesian analysis of HIC data Xu:2017obm , and is also lower than the data-driven result of the ALICE collaboration at $T\sim 150$ MeV ALICE:2021rxa . However, our results are in agreement with those obtained using the T-matrix formalism Tang:2023tkm ; Tang:2024dkz . In Fig. 19, we also show the latest results from the Catania group Sambataro:2024mkr obtained by using a refined version of their original quasi-particle description, known as QPMp. Their corresponding results are in agreement with our determination for bottom quark at low temperature, but remain much larger for charm quark and over-predict $D_{s}$ for both charm and bottom quarks for temperatures $T>250$ MeV.
@@ 6 Conclusions
In this paper, we present the first complete (2+1)-flavor lattice QCD calculation of the heavy-quark diffusion coefficient for a wide range of temperatures. For low temperatures, we extracted $\kappa$ at almost physical point, using a physical strange-quark mass $m_{s}$ and near-physical light-quark mass $m_{l}=m_{s}/20$, corresponding to a pion mass of $m_{\pi}\simeq 160$ MeV. On the other hand, for high temperatures we use a physical strange-quark mass $m_{s}$ and $m_{l}=m_{s}/5$, corresponding to a pion mass of $m_{\pi}\simeq 320$ MeV because light quark mass effects are not important in this region. This study extends previous calculations of the HotQCD collaboration Altenkort:2023oms ; Altenkort:2023eav to the physical point for low temperatures, extending the temperature range to 163 MeV $\leq T\leq$ 10000 MeV.
Our analysis of the chromo-electric correlation function employs the correct high-energy behavior of the corresponding spectral function, computed in Refs. delaCruz:2024cix ; Scheihing-Hitschfeld:2023tuz . Although this correction does not significantly change the extracted value of the heavy-quark diffusion coefficient within errors, it resolves the puzzle of the large rescaling factor $K_{E}$ observed in previous works Brambilla:2020siz ; Altenkort:2023oms . As discussed in Section 4.2, this improvement shows, for the first time, a direct agreement between lattice data and the perturbative NLO prediction at the level of the correlation function. This, in turn, reinforces the validity of our spectral function reconstruction strategy.
Our results show that close the crossover temperature, the spatial heavy-quark diffusion coefficient approaches the strongly coupled AdS/CFT estimate Casalderrey-Solana:2006fio ; Herzog:2006gh , supporting the picture of the quark-gluon plasma as a strongly-coupled system and an almost perfect fluid, in which heavy quarks rapidly thermalize Rapp:2018qla ; Dong:2019byy ; He:2022ywp . At high temperatures, our results are consistent within errors with the NLO perturbative QCD prediction Caron-Huot:2007rwy . Across the entire temperature range, the lattice QCD results remain systematically lower than the estimates from Bayesian reconstructions Xu:2017obm and data-driven extractions ALICE:2012ab , also showing a weaker dependence on the heavy-quark mass compared to certain phenomenological models.
@@ Appendix B Fits of the chromo-electric and chromo-magnetic correlators
Appendix C Quasi-particle model for charm and bottom degrees of freedom
@@
Figure 24: Left: Lattice results together with the interpolation of the charm susceptibility $\hat{\chi}_{2}^{C}$ from the BW collaboration Bellwied:2015lba with the recent $\hat{\chi}_{4}^{C}$ results from the HotQCD collaboration Kaczmarek:2025dqt . For comparison, we also show the prediction from a full hadron resonance gas model QM-HRG. Right: interpolation of the charm-quark mass extracted from the susceptibility data of the BW and HotQCD collaborations. The orange point indicates the interpolated value at $T=188$ MeV.
@@
Figure 25: Charm and bottom partial pressures in units of $T^{4}$. Left: charm quark partial pressure obtained by HotQCD determination Kaczmarek:2025dqt compared to the charm meson, charm baryon and total charm hadron pressure from a 1S1P HRG model. Right: bottom partial pressures, showing results from the 1S1P HRG (dotted lines) and full HRG (dashed lines). These estimates are subsequently used to extract the bottom-quark mass as a function of temperature across the different models discussed in the text.
@@
$T$ [MeV] $M_{c_{1}}$ ${\langle\mathbf{v}^{2}\rangle}_{c_{1}}$ $\langle\mathbf{\hat{p}}^{2}\rangle_{c_{1}}$ $M_{c_{2}}$ ${\langle\mathbf{v}^{2}\rangle}_{c_{2}}$ $\langle\mathbf{\hat{p}}^{2}\rangle_{c_{2}}$
@@
Table 3: Mean squared thermal velocity $\langle\mathbf{v}^{2}\rangle$} and scaled mean squared thermal momentum $\langle\mathbf{\hat{p}}^{2}\rangle\equiv{\langle\mathbf{p}^{2}\rangle}/(3M_{c}T)$ of charm degrees of freedom extracted from the quasi-particle models discussed in the text. Here, $M_{c}$ is the charm-quark mass in the corresponding quasi-particle model. All masses are given in GeV.
$T$ [MeV] $M_{b_{1}}$ ${\langle\mathbf{v}^{2}\rangle}_{b_{1}}$ $\langle\mathbf{\hat{p}}_{b_{1}}^{2}\rangle$ $M_{b_{2}}$ ${\langle\mathbf{v}^{2}\rangle}_{b_{2}}$ $\langle\mathbf{\hat{p}}_{b_{2}}^{2}\rangle$ $M_{b_{3}}$ ${\langle\mathbf{v}^{2}\rangle}_{b_{3}}$ $\langle\mathbf{\hat{p}}_{b_{3}}^{2}\rangle$
163 5.136 0.0852 1.1227 5.529 0.0824 1.0752 —— —— ——
174 5.104 0.0905 1.1364 5.482 0.0883 1.0811 —— —— ——
188 5.059 0.0972 1.1547 5.412 0.0958 1.0889 —— —— ——
195 5.059 0.1041 1.1141 —— —— —— 4.500 0.1174 1.1117
220 5.059 0.1168 1.1210 —— —— —— 4.500 0.1308 1.1265
251 5.059 0.1325 1.1285 —— —— —— 4.500 0.1470 1.1450
293 5.059 0.1519 1.1507 —— —— —— 4.500 0.1682 1.1702
352 5.059 0.1781 1.1824 —— —— —— 4.500 0.1966 1.2062
400 5.059 0.1984 1.2085 —— —— —— 4.500 0.2186 1.2358
444 5.059 0.2163 1.2326 —— —— —— 4.500 0.2380 1.2632
500 5.059 0.2383 1.2637 —— —— —— 4.500 0.2615 1.2985
1000 5.059 0.3982 1.5547 —— —— —— 4.500 0.4299 1.6304
Table 4: Mean squared thermal velocity $\langle\mathbf{v}^{2}\rangle$} and scaled mean squared thermal momentum $\langle\mathbf{\hat{p}}^{2}\rangle\equiv{\langle\mathbf{p}^{2}\rangle}/(3M_{b}T)$ of bottom degrees of freedom extracted from the quasi-particle models discussed in the text. Here, $M_{b}$ is the bottom-quark mass in the corresponding quasi-particle model. All masses are given in GeV.
@@
As discussed in the main text, a quasi-particle description is necessary to estimate the heavy-quark diffusion coefficient beyond LO in the inverse heavy-quark mass. In this section of the appendix, we detail the quasi-particle model used to compute $\langle\mathbf{v}^{2}\rangle$} and $\langle\mathbf{p}^{2}\rangle$}, based on available lattice QCD results. A model of this kind was first introduced in Ref. Petreczky:2008px to describe the temperature dependence of vector and axial-vector current susceptibilities for heavy (charm and bottom) quarks in quenched QCD. In Ref. Petreczky:2008px it was shown that $\langle\mathbf{v}^{2}\rangle$} obtained from the ratio of susceptibility of the vector current and the quark number susceptibility calculated on the lattice agrees with the quasi-particle model.
Generalized charm susceptibilities have been computed throughout the years in 2+1-flavor lattice QCD Bazavov:2014yba ; Bazavov:2023xzm ; Kaczmarek:2025dqt ; Bellwied:2015lba . Below the crossover temperature, their behavior can be understood within the framework of a hadron resonance gas (HRG) model. In this approach, the particle content includes not only experimentally observed charm hadrons listed by the Particle Data Group, but also additional states predicted by quark models Bazavov:2014yba ; Bazavov:2023xzm ; Kaczmarek:2025dqt . This extended version is known as the QM-HRG model.
Above the chiral crossover temperature, charm susceptibilities can be described using a model in which the total charm pressure is given by the sum of partial pressures from charm mesons, charm baryons, and charm quarks Mukherjee:2015mxc ; Bazavov:2023xzm ; Kaczmarek:2025dqt :